# What computer science has to say about the simulation hypothesis

### v4

@@ 4.1 Proof of the self-simulation lemma
Since the PCT holds for $V$, its evolution function $g(.,.,.)$ is computable (not just partial computable). Therefore if we fix $\Delta t$, the first argument of $g(.,.,.)$, the resultant dependence on the other two arguments is computable. If we now invoke the total recursion theorem (discussed in Appendix B), we see that there exists $n^{*}$ such that
@@ 4.2 Important features of the self-simulation lemma
It’s also worth pointing out that the proof of the self-simulation lemma does not require the full power of the recursion theorem. That theorem applies to any function as long as it is partial computable. However, the self-simulation lemma only needs to use it for the specific case of the evolution function, which is in fact a total computable function.
@@ 4.3 Difference between self-simulation lemma and previous work
It’s important to distinguish the self-simulation lemma from concepts considered in the earlier literature. First, the traditional versions of the simulation hypothesis discussed before did not involve the possibility that the beings running a simulating computer might be simulating their own universe, including themselves. Self-simulation is novel (pace informal musings like the one quoted in this paper’s epigraph).
Note as well that the self-simulation lemma is not a reformulation of the old warhorse, central to so much of computer science theory and the foundations of math, of the concept of a mathematical object “referencing itself”. To clarify the difference, specify some uniquely decodable function $f$ that takes two finite bit strings arguments. Consider any bit string $c$ in the range of $f$. So $f^{-1}(c)$ is the decoding of the bit string $c$ into an ordered pair of bit strings. Write those two bit strings as $f^{-1}_{1}(c)$ and $f^{-1}_{2}(c)$ respectively. So for all finite bit strings $a,b$,
$[f^{-1}_{1}(f(a,b)),f^{-1}_{2}(f(a,b))]=(a,b)$ (20)
Suppose we represent a universal TM $U$ as working on such encoded pairs of bit strings, by decoding them into a pair of bit strings, and then interpreting the first of those two bit strings as the specification of a TM which is to be run on the second of those two bit strings. So
$U(c)=[f^{-1}_{1}(c)][f^{-1}_{2}(c)]$ (21)
and therefore for all TMs $T$ and strings $w$ that we could feed into $T$,
$U(f(T,w))$ $=[f^{-1}_{1}(f(T,w))][f^{-1}_{2}(f(T,w))]$ (22)
$=T(w)$ (23)
Now suppose we choose $T$ in Eq. 23 to equal $U$ itself with $w$ arbitrary, as in the case of a UTM engaged in self-reference. Then if we apply Eq. 21 after having evaluated Eq. 23, we get
$U(f(U,w))$ $=U(w)$ (24)
$=[f^{-1}_{1}(w)][f^{-1}_{2}(w)]$ (25)
This is the output produced by whatever TM $f^{-1}_{1}(w)$ happens to be, if run on whatever input string $f^{-1}_{2}(w)$ happens to be. A priori, that output has no relation to the output of the UTM $U$ run on an encoding of itself. (Recall that $w$ is arbitrary.) Indeed, $f^{-1}_{1}(w)$ may not even halt when run on input string $f^{-1}_{2}(w)$.
@@
Indeed, for $U$ to simulate its actual behavior on its actual input string, we would need to feed $U$ an infinitely long input string, defined by an infinite regress of instances of itself
$f(f,f(f,(f,\ldots)))$ (26)
@@
(One can see this simply by iteratively expanding $w$ into $f(U,w)$.) It would be physically impossible for any physical computer $\mathcal{D}$ implementing $U$ to finish such a computation in finite (physical) time if it starts with this infinite string as its input. Moreover, there is no way to get information about the rest of the physical universe outside of that computer $\mathcal{D}$ into the input to $\mathcal{D}$ (since such information would need to be appended at the end of the infinite string $f(f,f(f,(f,\ldots)))$ None of these problems apply with the self-simulation lemma.
@@
It is also important to distinguish the self-simulation lemma from the age-old observation that if i) the universe extends infinitely spatially; ii) if the physical constants (and more generally physical laws) do not vary across that infinite space; iii) if the initial conditions of the regions inside the backward local light cones across the full universe include all physically possible initial conditions; then somewhere else in this universe there is a copy of each of us, identical down to a fine level of detail. This simple statistical phenomenon does not result in exact identity between any of those copies of us and ourselves. More importantly, perhaps the most striking aspect of the self-simulation lemma is that if it applies to us, it would mean that we are running a program on a computer that is simulating ourselves, as we run that simulating program. There is as direct a coupling between ourselves and our indistinguishable copy as possible; one of us is directly controlling the other one of us, but those two versions of us are one and the same. (See Section 6.) Indeed, if the self-simulation lemma holds, then there is an infinite regress of copies of you, residing inside the successive layers of nested dolls of computers simulating computers. [^13] See Section 5 for some of the mathematical properties that distinguish all of those nested simulations. Importantly, despite those distinguishing mathematical properties, if the Church-Turing thesis holds then any inhabitant of one of those nested simulations experiences the exact same reality as their copy in the other nested simulations. There is no experimental test that one can do to ascertain which nested simulation one inhabits, according to the Church Turing thesis. No such nested set of copies of an individual arises under the “age-old observation”.
@@
Another function considered in the literature that in some respects resembles the function $g$ of the self-simulation lemma is the Von Neumann constructor [68]. The Von Neumann constructor is a configuration of neighboring states in a particular cellular automaton that is capable of constructing an identical copy of itself as the cellular automaton evolves, resulting in two copies of the constructor. [^14] As an aside, the second recursion theorem of Kleene formally establishes how to design self-reproducing entities in a Turing machine. Despite Kleene’s renown, and the fact that he established this theorem many years before Von Neumann worked on his constructor, it seems that Von Neumann never emphasized Kleene’s precedence. However, a Von Neumann constructor releases the copy of itself once it has created it, and that copy then has a completely independent existence, undergoing different dynamics. In fact, there is no time at which the Von Neumann constructor and that of the copy of itself that it constructs undergo identical dynamics. The Von Neumann constructor does not “run” the copy of itself that it makes, in the sense of the function $g$ in the self-simulation lemma.
@@ 5.1 Necessity of time delay in self-simulation
One might suspect that there must be a delay between the time $\Delta t$ into the future that a self-simulating computer is simulating, and the time at which it completes that simulation. After all, suppose instead that ${\mathcal{T}}(\Delta t,w_{0},{\mathcal{S}}(\Delta t))=\Delta t$. This would imply that
@@
One could make a counter-argument though. In many ways our cosmological universe evolves more like a parallel computer rather than a serial one, so that for example the state space of $N$ could evolve like an infinite one-dimensional Cellular automaton. That would seem to allow an infinite number of operations to occur simultaneously — disproving the intuitive argument above. On the other hand, the computational implications if there were parallel rather than serial dynamics in our universe would result in some subtle problems. — see Appendix D. So this approach will not be considered here.
@@
Next, recall from Section 2.1 that in this paper I assume the encoding $\langle.,.\rangle$ produces strings whose lengths are non-decreasing functions of the lengths of its two arguments. So if $W$ has more than one state (and therefore $w_{\Delta t}$ is at least a bit long), $|\langle w_{\Delta t},n_{\Delta t}\rangle|>|n_{\Delta t}|$. In this case Eq. 36 is a contradiction. So our hypothesis must be wrong, which establishes the lemma for the case $|W|>1$. ∎
Lemma 3 means that for all $w_{0}$, ${\mathcal{T}}({\mathcal{S}}(\Delta t)),w_{0},\Delta t)$ has no upper bound as $\Delta t$ grows, and must always exceed $\Delta t$. It does not means that ${\mathcal{T}}(n^{*},w,\Delta t)$ is an everywhere increasing function of $\Delta t$ though; ${\mathcal{T}}({\mathcal{S}}(\Delta t),w_{0},\Delta t)+1$ can be less than ${\mathcal{T}}({\mathcal{S}}(\Delta t)),w_{0},\Delta t)$, as long as it’s greater than $\Delta t$.
One could of course forbid this possibility, simply by requiring that ${\mathcal{T}}({\mathcal{S}}(\Delta t)),w_{0},\Delta t)$ is not decreasing as a function of $\Delta t$. Another way to address this issue is to modify Definition 1 so that the simulating computer does not just produce a single future state of the universe being simulated, but rather produces an entire (finite) sequence of future states of the simulated universe, in order, as discussed in Section 2.2. A related way to address this issue is discussed in the next subsection.
@@ 5.2 Restrictions to impose on the evolution function
It will sometimes be natural to place several additional restrictions on the evolution function. Such restrictions do not change the main results presented below in Sections 3 and 4. However, it is helpful to invoke them in certain parts of the subsequent analysis of those main results.
@@
Fix $\Delta t$, and choose some integer $k$ such that $0<k<\Delta t$. A universe $V$ with a Markovian evolution function is shielded (for $\Delta t$ after iteration $k$) if for all $w,w_{0},n_{0}$,
@@
is independent of $w$, where as shorthand we defined
@@ 5.3 Cheating self-simulation and how to prevent it
The TM evolves shielded from its environment after this initialization. In all subsequent iterations up to $\Delta t$, $\underline{n}_{t}$ does not change, while $c$ increments by $1$ in each of those iterations. Then when the counter reaches $\Delta t$, it stops incrementing. At this time the contents of the tape of the TM is $(\Delta t,w_{0},{\Delta t})$, with the other variables of the TM (its state and its head’s position) having some value $r_{\Delta t}$. So the entire ID of the TM at this moment is ${\Delta t},w_{0},{\Delta t};r_{\Delta t}]$, and therefore the state of the full universe is $([\Delta t,w_{0},{\Delta t};r_{\Delta t}],w_{\Delta t})$.
Next the computer makes a second copy of $w_{0}$ and appends that together with $r_{\Delta t}$ to the end of its tape, so that it now has $\Delta t,w_{0},{\Delta t},r_{\Delta t},w_{0}]$ on its tape. [^16] Recall the discussion in Section 2.2 of the fact that it may prove convenient to augment our TMs with a special instruction that copies the contents of an arbitrarily large portion of $V$ to another portion of $V$ in a single iteration; that extra instruction could be used here, for calculational convenience, so that we don’t have to account for the number of iterations it would require a non-augmented TM to copy over that entire input, working on only a single variable in the multi-dimensional state space $W$ at a time. It then computes $w_{\Delta t}$ from that copy of $w_{0}$, so that at some later time $t_{2}>\Delta t$ its tape is $\Delta t,w_{0},{\Delta t},r_{\Delta t},w_{\Delta t}]$.
At this time the state of the full universe is $v_{t_{2}}=([\Delta t,w_{0},{\Delta t},r_{\Delta t},w_{\Delta t};r_{t_{2}}],w_{t_{2}})$. Note though that the value of the tape of the TM at $t_{2}$ is identical to the state of the entire universe at the earlier time, $\Delta t$. In other words, $n_{t_{2}}=g({\Delta t},w_{0},n_{0})$. We therefore satisfy the self-simulation lemma by choosing ${\mathcal{T}}(\Delta t,w_{0},n_{0})=t_{2}$.
One might argue that in a certain sense the algorithm used in Example 3 for self-simulation is a milder type of “cheating”, since the computer only increments a counter up to time $\Delta t$, not trying to evolve $w_{0}$ at all. However, we can nest that algorithm within itself an arbitrary number of times, so that $w_{0}$ is being evolved continually, not just after reaching the time $\Delta t$. This is illustrated in the following extension of Example 3, which shows how a shielded computer can produce a finite subsequence of an entire trajectory of states of the full universe, computing those states one after the other, in the proper time order:
@@
I refer to the procedure run in Example 4 as (greedy) nested simulation of a trajectory of states, with the sequence $\{t_{1},t_{2},\ldots\}$ called the simulation time sequence. [^17] The qualifier “greedy” reflects the fact that each successive computation is written to the tape of the TM as early as possible. Technically, greedy nested simulation means that for all $t_{i}$, $t_{i+1}-t_{i}$ is as small as possible. In the special case that nested simulation is applied by a universe to itself, as in Example 4, I refer to it (greedy) nested self-simulation as (greedy) nested self-simulation.
Note that in nested self-simulation, while the number of iterations to compute the state of the universe at a time $t$ is an increasing function of $t$, it is a partial function. Only a sub-sequence of the full trajectory of states of the universe defined by the simulation time sequence, $\{v_{t_{1}},v_{t_{2}},v_{t_{3}},\ldots\}$, is computed. This illustrates that the requirement of shielding does not totally prevent this particularly mild form of cheating. However, that requirement does restrict that kind of cheating to only work for a proper subset of all future times $t$. In other words, by requiring shielding we can ensure that only mild, “partial” cheating can occur.
As a final comment, note that nested self-simulation only places a sequence of pairs $(r_{t_{i}},w_{t_{i}})$ onto the tape, never separating the two elements of such a pair in its output. In addition, recall from above we can almost always treat $N$ as synonymous with $\underline{N}$, with Example 4 being the only instance in this paper in which we explicitly distinguish the $\underline{N}$ and $R$ components of $N$.
Given all this, one might think that we could absorb the variable $R$ into the variable $W$, leaving only $\underline{N}$ in the computer variable $N$. Doing that would change $W$ into the environment of $\underline{N}$, not of $N=\underline{N}\times R$. Using this alternative form of $W$ would have the advantage that it simplifies the notation of this paper.
However, Definition 6 requires that $R$ and $\underline{N}$ both lie in $N$. If we contradicted that definition by absorbing $R$ into $W$, then in nested self-simulation the computer $N$ (which would now only consist of the tape $\underline{N}$) would have to interact with its external environment for all iterations, never evolving autonomously. So we could not require that the computer be shielded from its environment. That’s why we don’t use this alternative form $W$ in this paper.

### v5

@@ 4.1 Proof of the self-simulation lemma
Since the PCT holds for $V$, its evolution function $g(.,.,.)$ is computable (not just partial computable). Therefore if we fix $\Delta t$, the first argument of $g(.,.,.)$, and then invoke the total recursion theorem, we see that there exists $n^{*}$ such that
@@ 4.2 Important features of the self-simulation lemma
It’s also worth pointing out that the proof of the self-simulation lemma does not require the full power of the recursion theorem. That theorem applies to any function so long as it is partial computable. However, the self-simulation lemma only needs to use it for the specific case of the evolution function, which is in fact a total computable function.
@@ 4.3 Difference between self-simulation lemma and previous work
It’s important to distinguish the self-simulation lemma from earlier concepts considered in the literature. First, the traditional versions of the simulation hypothesis discussed before did not involve the possibility that the beings running a simulating computer might be simulating their own universe, including themselves. Self-simulation appears to be novel (pace informal musings like the one quoted in this paper’s epigraph).
Note as well that the self-simulation lemma is not a reformulation of the old warhorse, central to so much of computer science theory and the foundations of math, of a mathematical object “referencing itself”. To clarify the difference, suppose we feed a universal TM $U$ an encoding of itself, $\langle U\rangle$, along with a finite bit string $w$ and then run that TM $U$. In other words, using $\langle a,b\rangle$ to indicate an encoding of a pair of bit strings, $a$, $b$, into a single bit string, we are considering the case where $U$’s input tape initially contains the bit string $\langle\langle U\rangle,w\rangle$. In this situation $U$ is “referencing itself”. In particular, $U(\langle\langle U\rangle,w\rangle)$ is computing the output bit string that it itself would calculate if its input tape were initialized with just $w$. Stated differently, the “inner” $U$, whose behavior if run on $w$ is being calculated, is the same as the “outer” $U$ running with $\langle\langle U\rangle,w\rangle$ on its input, with this same $w$.
There are several major differences between such self-reference and the self-simulation lemma. Most obviously, there is no notion of simulating the laws of physics arising in the computation $U(\langle\langle U\rangle,w\rangle)$. In contrast to the both the self-simulation lemma and the traditional version of the simulation hypothesis, neither the inner nor the outer $U$ is simulating the dynamics of our computational universe.
@@
Perhaps more importantly, $U(\langle\langle U\rangle,w\rangle)$ does not calculate what it itself would do if provided the string $\langle\langle U\rangle,w\rangle$ as input (rather than be provided $w$ as its input). So it is not simulating its actual behavior, but rather some counterfactual behavior. Indeed, for $U$ to simulate its actual behavior on its actual input string, we would need to feed $U$ an infinitely long bit string, defined by an infinite regress:
$\langle\langle U\rangle,\langle\langle U\rangle,\langle\langle U\rangle,\ldots\rangle\rangle\rangle\rangle\rangle\rangle)$ (20)
@@
(One can see this simply by iteratively expanding $w$ into $\langle\langle U\rangle,w\rangle)$.) It would be physically impossible for any computer $\mathcal{D}$ implementing $U$ to finish its computation in finite (physical) time if it starts with this infinite string as its input. Moreover, there is no way to get information about the rest of the physical universe outside of that computer $\mathcal{D}$ into the input to $\mathcal{D}$ (since such information would need to be appended at the end of the infinite string $\langle\langle U\rangle,\langle\langle U\rangle,\langle\langle U\rangle,\ldots\rangle\rangle\rangle\rangle\rangle\rangle)$). None of these problems apply with the self-simulation lemma.
@@
It is also important to distinguish the self-simulation lemma from the age-old observation that if i) the universe extends infinitely spatially; ii) the physical constants (and more generally physical laws) do not vary across that infinite space; iii) the initial conditions of the regions inside the backward local light cones across the full universe are IID random; then somewhere else in this universe there is a copy of each of us, identical down to a fine level of detail. [^9] In fact, as has often been pointed out, we don’t need to assume the IID random property. If the universe is infinite then the randomness of quantum mechanics means that there must be such a copy of each of us — indeed, there must be an infinite number of such copies.. This simple statistical phenomenon does not result in exact identity between any of those copies of us and us. More importantly, perhaps the most striking aspect of the self-simulation lemma is that if it applies to us, it would mean that we are running a program on a computer that is simulating ourselves, as we run that simulating program. There is as direct coupling between ourselves and our indistinguishable copy as possible; one of us is directly controlling the other one of us, but those two versions of us are one and the same. Phrased differently, one of of us is the parent, in a deep and fundamental sense, of the other. In fact, we are both the parent and the child. Indeed, if the self-simulation lemma holds, then there is an infinite regress of copies of you, residing inside the successive layers of nested dolls of computers simulating computers. No such nested set of copies of an individual arises under the “age-old observation”.
@@
Another function considered in the literature that in some respects resembles the function $g$ of the self-simulation lemma is the Von Neumann constructor. That is a configuration of neighboring states in a particular cellular automaton that is capable of constructing an identical copy of itself as the cellular automaton evolves. However, a Von Neumann constructor releases the copy of itself once it has created it, and that copy then has a completely independent existence, undergoing different dynamics. In fact, there is no time at which the Von Neumann constructor and that of the copy of itself that it constructs undergo identical dynamics. The Von Neumann constructor does not “run” the copy of itself that it makes, in the sense of the function $g$ in the self-simulation lemma.
@@ 5.1 Necessity of time delay in self-simulation
One might suspect that there must be a delay between the time $\Delta t$ into the future that a self-simulating computer is simulating, and the time at which it completes that simulation. Intuitively, the idea would be that, assume instead that ${\mathcal{T}}(\Delta t,w_{0},{\mathcal{S}}(\Delta t))=\Delta t$. This would imply that
@@
One could make a counter-argument though. One response to this intuitive argument is to point out that our cosmological universe in many ways evolves more like a parallel computer rather than a serial one, so that for example the state space of $N$ could evolve like an infinite one-dimensional Cellular automaton. That would see to allow an infinite number of operations to occur simultaneously — disproving the intuitive argument above. The computational implications of there being parallel rather than serial dynamics in our universe are subtle though — see Appendix D.
Even if we do restrict attention to a serial computer though, suppose we encode the bit-string $(w_{\Delta t},\langle w_{\Delta t},\langle w_{\Delta t},n_{\Delta t}\rangle\rangle),\ldots)$ as the output generated by a simple computer program, a program which would require very few bits to write down. So while $(w_{\Delta t},\langle w_{\Delta t},\langle w_{\Delta t},n_{\Delta t}\rangle\rangle),\ldots)$ is infinitely long, it can be encoded as a finite computer program (a short program, in fact). The simulation computer could write such a finite program onto its output tape in a finite time. On the other hand though, there is no computable function that can decode that version of the infinite string $(w_{\Delta t},\langle w_{\Delta t},\langle w_{\Delta t},n_{\Delta t}\rangle\rangle),\ldots)$ which has been encoded as a computer program. Any program that tries to do this would never halt.
@@
Next, recall that in this paper I assume the encoding $\langle.,.\rangle$ produces strings whose lengths are non-decreasing functions of the lengths of its two arguments. (See Section 2.1.) So if $W$ has more than one state (and therefore $w_{\Delta t}$ is at least a bit long), $|\langle w_{\Delta t},n_{\Delta t}\rangle|>|n_{\Delta t}|$. In this case Eq. 30 is a contradiction. So our hypothesis must be wrong, which establishes the lemma for the case $|W|>1$. ∎
Lemma 3 means that for all $w_{0}$, ${\mathcal{T}}({\mathcal{S}}(\Delta t)),w_{0},\Delta t)$ has no upper bound as $\Delta t$ grows, and must always exceed $\Delta t$. It does not means that ${\mathcal{T}}(n^{*},w,\Delta t)$ is an everywhere increasing function of $\Delta t$ though — ${\mathcal{T}}({\mathcal{S}}(\Delta t)),w_{0},\Delta t)+1$ can be less than ${\mathcal{T}}({\mathcal{S}}(\Delta t)),w_{0},\Delta t)$, so long as it’s greater than $\Delta t$.
One could of course forbid this possibility, simply by requiring that ${\mathcal{T}}({\mathcal{S}}(\Delta t)),w_{0},\Delta t)$ is not decreasing as a function of $\Delta t$. There are other ways to address this issue as well. One, mentioned in Section 2.2, is to modify Definition 1 so that the simulating computer does not just produce a single future state of the universe being simulated, but rather produces an entire (finite) sequence of future states of the simulated universe, in order. A related way to address this issue is addressed next, in Section 5.2.
@@ 5.2 Restrictions to impose on the evolution function
There are several additional restrictions it will sometimes be natural to place on the evolution function. It would not change the main results presented below in Sections 3 and 4 to impose those restrictions. However, it is helpful to invoke them in certain parts of the subsequent analysis of those results.
@@
Fix $\Delta t$, and choose some integer $k$ such that $0<k<\Delta t$. A universe $V$ with a Markovian evolution function is shielded (after iteration $k$) if for all $w,w_{0},n_{0}$,
@@
is independent of $w$, where we define
@@ 5.3 Cheating self-simulation and how to prevent it
The TM evolves shielded from its environment from now on. In all subsequent iterations up to $\Delta t$, $\underline{n}_{t}$ does not change, while $c$ increments by $1$ in each of those iterations. Then when the counter reaches $\Delta t$, it stops incrementing. At this time the contents of the tape of the TM is $(\Delta t,w_{0},{\Delta t})$, with the other variables of the TM (its state and its head’s position) having some value $r_{\Delta t}$. So the entire ID of the TM at this moment is ${\Delta t},w_{0},{\Delta t};r_{\Delta t}]$, and therefore the state of the full universe is $([\Delta t,w_{0},{\Delta t};r_{\Delta t}],w_{\Delta t})$.
Next the computer makes a second copy of $w_{0}$ and appends that together with $r_{\Delta t}$ to the end of its tape, so that it now has $(\Delta t,w_{0},{\Delta t},r_{\Delta t},w_{0})$ on its tape. [^11] Recall the discussion in Section 2.2 of the fact that it may prove convenient to augment our TMs with a special instruction that copies the contents of an arbitrarily large portion of $V$ to another portion of $V$ in a single iteration; that extra instruction could be used here, for calculational convenience, so that we don’t have to account for the number of iterations it would require a non-augmented TM to copy over that entire input, working on only a single variable in the multi-dimensional state space $W$ at a time. It then computes $w_{\Delta t}$ from that copy of $w_{0}$, so that at some later time $t_{2}>\Delta t$ its tape is $(\Delta t,w_{0},{\Delta t},r_{\Delta t},w_{\Delta t})$.
At this time the state of the full universe is $v_{t_{2}}=([\Delta t,w_{0},{\Delta t},r_{\Delta t},w_{\Delta t};r_{t_{2}}],w_{t_{2}})$. Note though that the value of the tape of the TM at $t_{2}$ is identical to the state of the entire universe at $\Delta t$. So $n_{t_{2}}=g({\Delta t},w_{0},n_{0})$. We therefore satisfy the self-simulation lemma by choosing ${\mathcal{T}}(\Delta t,w_{0},n_{0})=t_{2}$.
One might argue that in a certain sense the algorithm used in Example 3 for self-simulation is a milder type of “cheating”, since the computer only increments a counter up to time $\Delta t$, not trying to evolve $w_{0}$ at all. However, we can nest that algorithm within itself an arbitrary number of times, so that $w_{0}$ is being evolved continually, not just after reaching the time $\Delta t$. This is illustrated in the following extension of Example 3, which shows how a shielded computer can produce a subsequence of an entire trajectory of states of the full universe, computing those states one after the other, in the proper time order:
@@
I refer to the procedure run in Example 4 as (greedy) nested simulation of a trajectory of states, with the sequence $\{t_{1},t_{2},\ldots\}$ called the simulation time sequence. [^12] The qualifier “greedy” indicates the fact that each successive computation is written to the tape of the TM as early as possible. Technically, greedy nested simulation means that for all $t_{i}$, $t_{i+1}-t_{i}$ is as small as possible. In this case the nested simulation is applied by a universe to itself, a special case I refer to as greedy nested self-simulation. Note that in nested self-simulation, while the number of iterations to compute the state of the universe at a time $t$ is an increasing function of $t$, it is a partial function. Only a sub-sequence of the full trajectory of states of the universe defined by the simulation time sequence, $\{v_{t_{1}},v_{t_{2}},v_{t_{3}},\ldots\}$, is computed.
The nested self-simulation only places a sequence of pairs $(r_{t_{i}},w_{t_{i}})$ onto the tape, never separating the two elements of such a pair in its output. In addition, recall from above we can almost always treat $N$ as synonymous with $\underline{N}$, with Example 4 being the only instance in this paper in which we explicitly distinguish the $\underline{N}$ and $R$ components of $N$. Given all this, one might think that formally, we could absorb the variable $R$ into the variable $W$, leaving only $\underline{N}$ in the computer variable $N$. Doing that would change $W$ into the environment of $\underline{N}$, not of $N=\underline{N}\times R$.
This would simplify the notation of this paper. However, Definition 6 requires that $R$ and $\underline{N}$ both lie in $N$. If we absorbed $R$ into $W$, then in nested self-simulation the computer $N$ (which would now only consist of the tape $\underline{N}$) would have to interact with its external environment for all iterations, never evolving autonomously. So we could not require that the computer be shielded from its environment.